B) Oral Drug Medication Oral medication, in tablet or capsule form, will have a dosage strength per unit written on the label. For example: Penicillin (250 mg. per tablet). Many times the dosage prescribed does not match the dosage within individual tablets or capsules. When giving oral medication, tablets are often scored or grooved so that an accurate dosage can be given when the prescription dose is lower than the available medication. A 10% difference between prescription dosage and dosage available is acceptable for most applications. Below are some examples of the difficulties in oral medication determination. Calculate the correct dosage to be taken in each case and please show your work 1. The physician prescribes Lasix, 80 miligrams p.o. stat. The pharmacy has packages of 20 milligrams Lasix tablets available. How many tablets are necessary? 2.The physician prescribes 75 mg of Tylenol elixir p.o. q.4h. p.r.n. for pain. The elixir is available in 25 mg/5ml. How many teaspoons will you give? (Note: one teaspoon is 5 ml) 3. The physician prescription is for 0.30 g of Tegretol p.o. ti.d. The Tegretol is available from the pharma cy in a suspension of 100 mg/5ml. How many teaspoons will you give? 53
Answer #1

1.) As per prescription 80 mg lasix is to be taken. Tablets available in the pharmacy are of 20 mg.

` 1 tablet = 20 mg

4 tablets = 4x 20 mg = 80 mg

Therefore 4 tablets are necessary.

2.) Prescribed dosage of tylenol is 75 mg.

Available tylenol is 25 mg / 5 ml and 1 teaspoon is 5 ml

1 teaspoon =25 mg

25mg = 1 teaspoon

1mg = 1/25 tsp

75 mg = 75/25 tsp

75 mg =3 tsp

Therefore 3 teaspoons of 25 mg/5ml tylenol should be given.

3.) Prescribed dosage of tegretol is 0.30g or 300mg

1g =1000mg

0.3 g = 0,3 x 1000 = 300 mg

In pharmacy 100mg/5 ml tegretol is available.

1 teaspoon is 5 ml

1 tsp =100 mg

100 mg =1 tsp

1mg = 1/100 tsp

300mg = 300/100 tsp =3 tsp

3 teaspoon of 100mg/5 ml tegretol should be given